OpenStudy (anonymous):

Consider this equation: 4x – 5 + 6x = 6 Which is an equivalent equation after combining like terms? A. -1 + 6x = 6 B. 10x - 5 = 6 C. 4x + 1 = 6 D. 4x - 5 =

OpenStudy (godlovesme):

Have u tried to combine the like terms? :)

OpenStudy (yanasidlinskiy):

In other words, do you know what like terms are?

OpenStudy (anonymous):

Yea i got c but idk how i feel about it

OpenStudy (yanasidlinskiy):

What do you get when you add 4x+5x?

OpenStudy (anonymous):

9x

OpenStudy (yanasidlinskiy):

I meant 6x. Not 5x.

OpenStudy (anonymous):

10x

OpenStudy (yanasidlinskiy):

So now you're left with: 10x-5=6

OpenStudy (anonymous):

so add 5?

OpenStudy (yanasidlinskiy):

Yes, you'd add 5 to solve the rest of the equation. But they're only asking you only to combine terms. And that is what you get when you combine terms.

OpenStudy (yanasidlinskiy):

Combined terms are supposed to be on one side of the equation for example: 5x+3x+ 2+10=25 To solve we would just add the like terms. 8x+12=25. Do you see where I'm going with this?

OpenStudy (anonymous):

yes

OpenStudy (yanasidlinskiy):

So makes sense, correct? There you go! You do know what's the final answer, right?

OpenStudy (anonymous):

no still lost im gonna guess its b

OpenStudy (yanasidlinskiy):

How are you lost? I just showed you the steps.